# Exposure Completing for Temporally Consistent Neural High Dynamic Range Video Rendering

### v1

@@ Abstract.
High dynamic range (HDR) video rendering from low dynamic range (LDR) videos where frames are of alternate exposure encounters significant challenges, due to the exposure change and absence at each time stamp. The exposure change and absence make existing methods generate flickering HDR results. In this paper, we propose a novel paradigm to render HDR frames via completing the absent exposure information, hence the exposure information is complete and consistent. Our approach involves interpolating neighbor LDR frames in the time dimension to reconstruct LDR frames for the absent exposures. Combining the interpolated and given LDR frames, the complete set of exposure information is available at each time stamp. This benefits the fusing process for HDR results, reducing noise and ghosting artifacts therefore improving temporal consistency. Extensive experimental evaluations on standard benchmarks demonstrate that our method achieves state-of-the-art performance, highlighting the importance of absent exposure completing in HDR video rendering. The code is available at https://github.com/cuijiahao666/NECHDR.
@@
copyright: acmlicensedjournalyear: 2024doi: XXXXXXX.XXXXXXXconference: Make sure to enter the correct conference title from your rights confirmation email; October 28 - November 1, 2024; Melbourne, Australia.isbn: 978-1-4503-XXXX-X/18/06ccs: Computing methodologies Artificial intelligenceccs: Computing methodologies Computer visionccs: Computing methodologies Image and video acquisitionccs: Computing methodologies Computational photography
@@ 1. Introduction
In the proposed NECHDR framework, pyramid features of input LDR frames are extracted by the feature encoder, then fed into the exposure completing decoder and the HDR rendering decoder. The exposure completing decoder interpolates the features of neighbor LDR frames at every level of the feature pyramid. The interpolated LDR features are combined with the features from input LDR images as the inputs to HDR rendering decoder. The HDR rendering decoder estimates coarse HDR results and the optical flows at every level. The flows can facilitate feature interpolating in exposure completing decoder.
At the end of NECHDR, a simple blending network is used to integrate interpolated LDR frames, input frames, and coarse HDR frames, which can achieve high-quality and temporally consistent HDR reconstruction results.
@@ 2.1. HDR Image Rendering
HDR imaging technology aims to extend the dynamic range of images by merging LDR images with different exposures. Both the HDR image and video tasks encounter the misalignment between LDR images. Pioneer studies for HDR image rendering employ image alignment techniques to address this issue. Early works (Oh et al., 2014; Khan et al., 2006; Heo et al., 2010; Zhang and Cham, 2011; Lee et al., 2014; Grosch et al., 2006) often align LDR images globally, then detect and discard unaligned regions.
@@ 3.2. Exposure Completing
Our exposure completing process takes the neighbor features $\{\Phi_{t-1},\Phi_{t+1}\}$ and the optical flows $\{F_{t-1\rightarrow t},F_{t+1\rightarrow t}\}$ as input. During this process for the features with absent exposure information, the $k$-th level neighbor features $\{\phi_{t-1}^{k},\phi_{t+1}^{k}\}$ and optical flows $\{f_{t-1\rightarrow t}^{k},f_{t+1\rightarrow t}^{k}\}$ are processed to obtain completed feature $\widehat{\phi}_{t}^{k}$:

### v2

@@ Abstract.
High dynamic range (HDR) video rendering from low dynamic range (LDR) videos where frames are of alternate exposures encounters significant challenges, due to the exposure change and absence at each time stamp. The exposure change and absence make existing methods generate flickering HDR results. In this paper, we propose a novel paradigm to render HDR frames via completing the absent exposure information, hence the exposure information is complete and consistent. Our approach involves interpolating neighbor LDR frames in the time dimension to reconstruct LDR frames for the absent exposures. Combining the interpolated and given LDR frames, the complete set of exposure information is available at each time stamp. This benefits the fusing process for HDR results, reducing noise and ghosting artifacts therefore improving temporal consistency. Extensive experimental evaluations on standard benchmarks demonstrate that our method achieves state-of-the-art performance, highlighting the importance of absent exposure completing in HDR video rendering. The code is available at https://github.com/cuijiahao666/NECHDR.
@@
copyright: acmlicensedjournalyear: 2024copyright: acmlicensedconference: Proceedings of the 32nd ACM International Conference on Multimedia; October 28-November 1, 2024; Melbourne, VIC, Australiabooktitle: Proceedings of the 32nd ACM International Conference on Multimedia (MM ’24), October 28-November 1, 2024, Melbourne, VIC, Australiadoi: 10.1145/3664647.3680935isbn: 979-8-4007-0686-8/24/10ccs: Computing methodologies Artificial intelligenceccs: Computing methodologies Computer visionccs: Computing methodologies Image and video acquisitionccs: Computing methodologies Computational photography
@@ 1. Introduction
In the proposed NECHDR framework, pyramid features of input LDR frames are extracted by the feature encoder, then fed into the exposure completing decoder and the HDR rendering decoder. The exposure completing decoder interpolates the features of neighbor LDR frames at every level of the feature pyramid. The interpolated LDR features are combined with the features from input LDR images as the inputs to HDR rendering decoder. The HDR rendering decoder estimates coarse HDR results and the optical flows at every level. The flows can facilitate feature interpolating in exposure completing decoder. At the end of NECHDR, a simple blending network is used to integrate interpolated LDR frames, input frames, and coarse HDR frames, which can achieve high-quality and temporally consistent HDR reconstruction results.
@@ 2.1. HDR Image Rendering
HDR imaging technology aims to extend the dynamic range of images by merging LDR images with different exposures. Both the HDR image and video reconstructing tasks encounter the misalignment between LDR images. Pioneer studies for HDR image rendering employ image alignment techniques to address this issue. Early works (Oh et al., 2014; Khan et al., 2006; Heo et al., 2010; Zhang and Cham, 2011; Lee et al., 2014; Grosch et al., 2006) often align LDR images globally, then detect and discard unaligned regions.
@@ 3.2. Exposure Completing
Our exposure completing process takes the neighbor features $\{\Phi_{t-1},\Phi_{t+1}\}$ and the optical flows $\{F_{t-1\rightarrow t},F_{t+1\rightarrow t}\}$ as input. During this process, for the features with absent exposure information, the $k$-th level neighbor features $\{\phi_{t-1}^{k},\phi_{t+1}^{k}\}$ and optical flows $\{f_{t-1\rightarrow t}^{k},f_{t+1\rightarrow t}^{k}\}$ are processed to obtain completed feature $\widehat{\phi}_{t}^{k}$:
@@ 5. Conclusion
Appendix A More quantitative results
In Table 4 and Table 5, we provide the results of the current and proposed methods on the more recent metrics under the two-exposure setting, which are all tested on the CinematicVideo dataset (Froehlich et al., 2014). As shown in the Table 4, our method has achieved the state-of-the-art performance on the more recent image quality metrics. We also conduct comparisons on the video quality metric ForVideoVDP in Table 5, which considers information in temporal domain. Our method achieved state-of-the-art performance on this metric also.
Table 4. Additional results on more recent metrics
[…]
Chen 9.22 34.24 0.9190 34.45 0.9188
LAN-HDR 9.05 32.30 0.9017 33.11 0.9041
HDRFlow 9.28 33.24 0.9146 33.47 0.9168
Ours 9.30 34.66 0.9292 34.84 0.9290
Table 5. The performance on video metrics: ForVideoVDP
Methods Chen LAN-HDR HDRFlow Ours
ForVideoVDP 8.9950 8.3404 8.8692 9.0680
Appendix B The efficiency of the proposed methods
In Table 6, we provide a comparison of the efficiency between our proposed method and the current state-of-the-art methods on the CinematicVideo dataset (Froehlich et al., 2014) under the two-exposure setting. In both settings, we offer lightweight versions of the NECHDR network by reducing the number of feature channels. In the lightweight versions, feature channels from the first to the fourth levels in feature encoder $\mathcal{E}$ is set to 24, 36, 54, 72, respectively. The channels of both the exposure completing decoder $\mathcal{D}_{I}$ and HDR rendering decoder $\mathcal{D}_{R}$ are adjusted accordingly. Furthermore, the number of feature channels in the third and fifth convolution layers of HDR rendering decoder $\mathcal{D}_{R}$ will be adjusted to 24. As shown in Table 7, reducing the number of feature channels significantly decreases the parameter and inference time of our model, with no noticeable change in performance.
Table 6. The efficiency of current and proposed methods
Methods Flops(G) Para.(M) Time(ms) PSNRT SSIMT
Chen - 6.1 522 35.65 0.8949
LAN-HDR 3046 7.3 415 38.22 0.9100
HDRFlow 367 3.3 55 39.30 0.9156
Ours 1376 8.8 178 40.59 0.9241
Table 7. Quantitative comparisons with the lightweight version methods “NECHDR-lw”.
two-Exposures three-Exposures
NECHDR NECHDR-lw NECHDR NECHDR-lw
Para. (M) 8.8 4.2 ($\downarrow 52\%$) 9.9 4.8 ($\downarrow 52\%$)
Time (s) 0.178 0.138($\downarrow 23\%$) 0.163 0.130 ($\downarrow 20\%$)
PSNRT 40.59 40.33 ($\downarrow 0.7\%$) 37.24 37.11 ($\downarrow 0.3\%$)
SSIMT 0.9241 0.9217 ($\downarrow 0.3\%$) 0.9102 0.9053 ($\downarrow 0.5\%$)
Appendix C The ablation study of loss functions.
The ablation study of the loss functions is provided in Table 8. A significant performance improvement occurs after adding the exposure completion loss $\mathcal{L}^{com}_{I}$, which validates our core contribution: the importance of exposure completion for HDR video rendering. The losses for features are introduced in  (Kong et al., 2022) to provide useful low-level structure information. Here, the losses can enhance the details of HDR rendering, thereby improving performance. Finally, combined with optical flow loss $\mathcal{L}_{F}$, our method achieves the final performance.
Table 8. The ablation of losses in our method
$\mathcal{L}^{ren}_{I}$ $\mathcal{L}^{com}_{I}$ $\mathcal{L}^{ren}_{G}$ $\mathcal{L}^{com}_{G}$ $\mathcal{L}_{F}$ PSNRT SSIMT
39.83 0.9213
40.46 0.9225
40.50 0.9231
40.52 0.9932
40.59 0.9941
Figure 9. Details of the feature encoder $\mathcal{E}$
Figure 10. Details of the exposure completing decoder $\mathcal{D}_{I}$
Figure 11. Details of the HDR rendering decoder $\mathcal{D}_{R}^{1}$. Left: two-exposure; right: three-exposure.
Figure 12. Details of the HDR rendering decoder $\mathcal{D}_{R}^{2}$.
Figure 13. Details of the HDR rendering decoder $\mathcal{D}_{R}^{3}$.
Figure 14. Details of the HDR rendering decoder $\mathcal{D}_{R}^{4}$.
Appendix D More Qualitative comparisons
Under the three-exposure setting, we provide more qualitative comparison results in Fig. 7 and Fig. 8 between our method and the current state-of-the-art approaches: Chen (Chen et al., 2021), LAN-HDR (Chung and Cho, 2023), HDRFlow (Xu et al., 2024). As shown in Fig. 7, we obtain the completed frames “EC1” and “EC2”. The “EC1” provides high-exposure information with high signal-to-noise ratio to assist in removing noise from the dark background of the reference frame, while “EC2” offers low-exposure foreground object details to restore details missing due to overexposure in the reference frame. Therefore, our proposed method achieves the HDR rendering result closest to ground truth. We also provide two scenarios in Fig. 8, where motions occur with saturation and noise respectively. In the left image of Fig. 8, although “EC1” is not well completed due to the large noise and motion of the first and fourth input frames, the well-completed “EC2” provides information about the saturated regions for HDR rendering, resulting in ghost-free HDR result. In the right image of Fig. 8, the completed frame “EC2” provides detailed information with low noise, thus achieving the lowest noise level in these methods. The above visualization results elucidate the performance enhancement of our method under the three-exposure setting, further proving the robustness of our approach across different exposure settings.
Appendix E Network Architecture
We provide details of the NECHDR network we proposed, mainly including the following sub-networks: feature encoder $\mathcal{E}$, exposure completing decoder $\mathcal{D}_{I}$, $k$-levels HDR rendering decoders $\{\mathcal{D}_{R}^{k}|k=1,2,3,4\}$ and blending network. Both The feature encoder $\mathcal{E}$ and exposure-completing decoder $\mathcal{D}_{I}$ are weight-shared in the two-exposure setting or the three-exposure setting. We take input LDR frames with a size of $256*256$ as example and illustrate in figures for above sub-networks. In Fig. 91011121314, the parameters for “Conv” and “Deconv”, listed from left to right, are input channels, output channels, kernel size, stride, and padding. In our network, each “Conv” is followed by a PReLU , whereas there is no activation subsequent to each ’Deconv’ layer.
In the two-exposure setting, the parameter shared feature encoder $\mathcal{E}$ in Fig. 9 outputs pyramid features for the three input LDR frames $\{l_{t-1},l_{t},l_{t+1}\}$, while for the five input LDR frames $\{l_{t-2},l_{t-1},l_{t},l_{t+1},l_{t+2}\}$ in the three-exposure setting. As for the exposure completing decoder $\mathcal{D}_{I}$ in the Fig. 10, we adopt the same structure and parameters in both settings. Particularly, in the three-exposure setting, the parameter shared decoder $\mathcal{D}_{I}$ is utilized twice, producing two interpolated frames $\{\widehat{l}_{t}^{\epsilon_{1}},\widehat{l}_{t}^{\epsilon_{2}}\}$ and theirs corresponding features $\widehat{\Phi}_{t,1}^{k}=\{\widehat{\phi}_{t,1}^{k}|k=1,2,3\},\widehat{\Phi}_{t,2}^{k}=\{\widehat{\phi}_{t,2}^{k}|k=1,2,3\}$. The details of HDR rendering decoder $\mathcal{D}_{I}^{k}$ in $k$ levels, where $k=1,2,3,4$, are shown in Fig. 11121314, respectively. In each of the Fig. 11121314, the left illustrates the $k$-th level $\mathcal{D}_{I}^{k}$ under two-exposure setting, while the right is under three-exposure setting. The details regarding the blending network can be found in (Xu et al., 2024).